The weekly wages of employees of volta gold normally distributed about a mean of $1250 with a standard deviation of $120. Find the probability of an employee having a weekly weekly wage lying. I. Between $1,320 and $970. ii. Under $1,400. iii. Over $1,290
1
Expert's answer
2021-07-19T07:29:59-0400

"\mu=1250 \\
\sigma=120 \\
i. \; P(970<X<1320) = P(X<1320) -P(X<970) \\
= P(Z< \frac{1320-1250}{120}) -P(Z< \frac{970-1250}{120}) \\
= P(Z< 0.5833) -P(Z< -2.333) \\
= 0.72005-0.00982 \\
= 0.71023 \\
ii. \;P(X<1400) = P(Z< \frac{1400-1250}{120}) \\
=P(Z< 1.25) \\
= 0.8943 \\
iii. \; P(X>1290) = 1 -P(X<1290) \\
= 1 -P(Z< \frac{1290-1250}{120}) \\
= 1 -P(Z<0.3333) \\
= 1 -0.63043 \\
= 0.36957"